OpenStudy (anonymous):

in triangle abc, angle c=120degress cos square a + cos square b + cosa.cosb=?

OpenStudy (shubhamsrg):

Is there any other info given ? About the triangle or anything ?

OpenStudy (anonymous):

no... angle c is 120 degree and we need tofind find \[\cos ^{2}a+\cos ^{2}b+cosa.cosb\] options are: a)3/4 b) 3/2 c) 1/2 d) 1/4

OpenStudy (shubhamsrg):

@ganeshie8

OpenStudy (anonymous):

wat happened

OpenStudy (shubhamsrg):

I am unable to solve this. But I tagged @ganeshie8 . He'll probably be able to help you.

OpenStudy (anonymous):

okay thank you

OpenStudy (shubhamsrg):

Nah, I don't think that is correct. a=b=30 degrees does not yield any of the options.

OpenStudy (anonymous):

the quesiotn and options both are correct....a=b=30 is not necessary... take a+b=60 and try to split th multiplication xpressions into additive expressions... eg. 2cos((C+D)/2)cos((C-D)/2)=cos(C+D)+cos(C-D)

OpenStudy (anonymous):

im not able to get it

ganeshie8 (ganeshie8):

try below : A+B = 60 c^2 = a^2 + b^2 + 1 cos^2A + cos^2B + cosAcosB = (cosA + cosB)^2 - cosAcosB

ganeshie8 (ganeshie8):

I think @shubhamsrg is right, the options are incorrect as a=b=30 does not yield any of the options